### v2

# Genesis-Metallicity: Universal Non-Parametric Gas-Phase Metallicity Estimation

Abstract
We introduce genesis-metallicity, a gas-phase metallicity measurement python software employing the direct and strong-line methods depending on the available oxygen lines. The non-parametric strong-line estimator is calibrated based on a kernel density estimate in the 4-dimensional space of O2 = [O II]/H; O3 = [O III]/H; H equivalent width EW(H); and gas-phase metallicity . We use a calibration sample of 1510 galaxies at with direct-method metallicity measurements, compiled from the JWST/NIRSpec and ground-based observations. In particular, we report 122 new NIRSpec direct-method metallicity measurements at . We show that the O2, O3, and EW(H) measurements are sufficient for a gas-phase metallicity estimate that is more accurate than 0.09 dex. Our calibration is universal, meaning that its accuracy does not depend on the target redshift. Furthermore, the direct-method module employs a non-parametric (O II) electron temperature estimator based on a kernel density estimate in the 5-dimensional space of O2, O3, EW(H), (O II), and (O III). This (O II) estimator is calibrated based on 1004 spectra with detections of both [O III] and [O II], notably reporting 20 new NIRSpec detections of the [O II] doublet. We make genesis-metallicity and its calibration data publicly available and commit to keeping both up-to-date in light of the incoming data.
1 Introduction
The “direct-method” provides a highly reliable measure of the gas-phase metallicity in galaxies. However, this method relies on an estimate of the electron temperature before the ionic abundances can be derived from the abundance-sensitive emission lines. Unfortunately, the often-faint temperature-sensitive emission lines such as [O III] and [O II] remain mostly elusive in large spectroscopic surveys, hindering the application of the direct-method on large samples. This has made the “strong” emission lines such as the [O II] and [O III] doublets the most commonly used proxies for the gas-phase metallicities of galaxies with available rest-optical spectroscopy (see, e.g., Savaglio et al., 2005; Erb et al., 2006; Maiolino et al., 2008; Mannucci et al., 2009; Zahid et al., 2011, 2014; Wuyts et al., 2012, 2016; Belli et al., 2013; Henry et al., 2013; Kulas et al., 2013; Cullen et al., 2014; Yabe et al., 2014; Maier et al., 2014; Steidel et al., 2014; Troncoso et al., 2014; Kacprzak et al., 2016, 2015; Sanders et al., 2015, 2021; Hunt et al., 2016; Onodera et al., 2016; Suzuki et al., 2017; Curti et al., 2017, 2024a; Langeroodi et al., 2023; Langeroodi & Hjorth, 2023; Heintz et al., 2023; Nakajima et al., 2023; Chemerynska et al., 2024; Sarkar et al., 2024). This practice is commonly known as the “strong-line” metallicity estimation: several polynomial relations between various strong-line ratios and gas-phase metallicity are calibrated either empirically on samples with direct-method measurements (see, e.g., McGaugh, 1991; Pilyugin et al., 2010; Pilyugin & Grebel, 2016; Curti et al., 2017; Jiang et al., 2019; Nakajima et al., 2022) or against the predictions of photoionization models (see, e.g., McCall et al., 1985; Denicoló et al., 2002; Kewley & Dopita, 2002; Hirschmann et al., 2023).
Despite the success of traditional strong-line metallicity estimators in enabling statistically significant chemical enrichment studies across a wide range of galaxy properties and redshift (see references above), they come with nuanced caveats rooted in their “parametric” nature. Firstly; the 2D projections of the calibration data onto the line ratio vs. metallicity planes risk overlooking the complexities of the higher-order parameter space. Even the 2D projections are often too complex to be fully captured by polynomials. For instance, particularly at low metallicities, large scatter is reported around the best-fit O2– and O32– relations. Nakajima et al. (2022) showed that the offsets from these best-fit relations depend on the ionization state of interstellar media (ISM), and can be captured by the equivalent width of H, EW(H).
Second; the parametric calibrations are prone to “hot” spots which render the estimates in certain metallicity windows highly uncertain. For instance, the best-fit polynomials to the O3– and R23– projections are widely used as primary metallicity estimators because these relations exhibit relatively tight scatter. However, both projections are non-monotonic, with a turnover metallicity of . This means that i) multiple metallicity solutions exist for each input O3 and R23, which should be sifted based on other projections; ii) the metallicity estimation around the turnover value is highly uncertain due to the flattening of the calibration curve; and iii) observed line ratios higher than the maximum allowed by the calibration curve universally yield the turnover metallicity, failing to capture the intrinsic scatter of the relation.
Third; recent parametric calibrations at high redshifts based on NIRSpec spectroscopy indicate noticeable deviations from the local-universe calibrations (Sanders et al., 2024; Laseter et al., 2024), potentially suggesting a non-universality in the strong-line method. However, as shown by Nakajima et al. (2018) and Nakajima et al. (2022), the 2D-projected relationships between the line ratios and gas-phase metallicity are influenced by the ionization parameter. Therefore, high-redshift deviations from the locally-calibrated parametric strong-line estimators are expected, as the high-redshift galaxies exhibit systematically higher ionization parameters. This is evidenced by their observed extremely high O32, EW(H), and EW(H) values (Langeroodi et al., 2023; Langeroodi & Hjorth, 2024; Rinaldi et al., 2023), indicative of high ionization parameters (Kewley & Dopita, 2002; Hirschmann et al., 2023) and bursty star formation histories (Smit et al., 2016; Langeroodi & Hjorth, 2024). Nonetheless, it is essential for any strong-line calibration to capture such dependencies and remain insensitive to these systematics.
Here, we overcome these caveats by developing a “non-parametric” strong-line metallicity estimator. We achieve this by a kernel density estimate (KDE; Silverman, 1986; Scott, 1992) of the probability density function (PDF) in the multi-dimensional space of emission line observables and gas-phase metallicity (Section 4). This PDF is then used to estimate the gas-phase metallicity for any combination of input emission line observables. We calibrate our strong-line estimator on a sample of 1510 galaxies at with direct-method metallicity measurements, the largest of such compilations to date (Sections 2 and 3). In particular, we report 122 new direct-method metallicity measurements at based on NIRSpec multi-shutter assembly (MSA; Jakobsen et al., 2022; Ferruit et al., 2022) spectroscopy; this corresponds to a fold increase in the sample size of directly-measured metallicities. We show that the O2, O3, and EW(H) measurements are sufficient for a gas-phase metallicity estimate that is more accurate than 0.09 dex. Our calibration is universal, meaning that its accuracy does not depend on the target redshift. We make genesis-metallicity (Langeroodi, 2025) and its calibration data available at https://github.com/langeroodi/genesis_metallicity (the calibration data will be made available upon publication).
2 Data
In this Section, we present an overview of the spectra utilized in our strong-line metallicity calibration. This data consists of 1510 spectra with direct-method metallicity measurements, including 171 galaxies observed with the NIRSpec MSA, 122 of which are reported for the first time in this work and the rest are taken from the literature (Section 2.1); 1213 galaxies observed with ground-based instruments (Section 2.2); and 126 high-metallicity spectra generated by stacking the SDSS spectra (Section 2.3). Figure 1 provides an overview of this sample. The line fluxes are reported in Table 1. We note that the H, H, H, and H Balmer lines are used for dust reddening correction of emission lines. For this purpose, we assumed a Calzetti et al. (2000) dust curve111Several recent studies have found evidence that while non-negligible dust is already present in early-universe galaxies, their attenuation curves might deviate from those in local universe (Langeroodi et al., 2024; Ciesla et al., 2024; Burgarella et al., 2024, 2025; Markov et al., 2025a, b; McKinney et al., 2025; Fisher et al., 2025). As such, a dedicated investigation of potential systematic biases caused by adopting the Calzetti et al. (2000) or other similar dust curves in this and other similar work is timely. However, this goes beyond the scope of our current study. and case-B recombination222The dust attenuation module is available at https://github.com/langeroodi/genesis_metallicity.
2.1 NIRSpec
We searched the JADES DR3 (D’Eugenio et al., 2024) NIRSpec MSA medium-resolution333Due to its low spectral resolution, the prism grating almost never resolves the [O III] line from H. Exceptions can occur at , where these lines fall at relatively high-resolution () prism wavelengths (Williams et al., 2023; Schaerer et al., 2024; Curti et al., 2024b). spectra for [O III] and [O II] detections. For this purpose, we used pPXF (Cappellari & Emsellem, 2004; Cappellari, 2017, 2022) to measure the emission line fluxes. For the objects covered in multiple JADES observations, we stacked the spectra from repeated gratings to enhance the signal. We adopted the spectroscopic redshifts reported by the JADES team as a starting point, and for each object ran pPXF on the medium-resolution spectra covering its [O III] and [O II] emission. We then visually inspected the subsample with either [O III] or [O II] flux signal-to-noise ratios (S/N) greater than 3. We confirm 138 galaxies with robust [O III] detections (), 20 of which also exhibit robust [O II] detection (). These exclude the confirmed broad-line AGN from Maiolino et al. (2024). We also fitted the prism spectra of the [O III]-detected sample to achieve full coverage of the [O II]; H; [O III]; H; [O III]; H; and [O II] lines. The NIRSpec medium-resolution spectra as well as the corresponding best-fit pPXF models for some example galaxies with [O III] and [O II] detections are presented in Appendix A and Figures 7, 8, 9, 10, 11, 12, 13, and 14.
We combined the medium-resolution and prism line flux measurements into a final catalog. We exclusively used the medium resolution measurements for the H, [O III], and [O II] lines. This is because in prism spectra H and [O III] lines are rarely deblended from one another and [O II] often appears too faint to be confidently distinguished from the continuum. Since the H to H flux ratio is the highest-signal Balmer line ratio used to correct for dust attenuation, we prioritized measuring both on the same grating to avoid cross-grating calibration offsets. If multiple gratings provided simultaneous high-significance detections () of both lines, we prioritized medium-resolution gratings as they generally resolve H from [O III] much more comfortably. The EW(H) is calculated using the pPXF best-fit continuum on the same grating where the H flux is read. For the rest of the lines, if high-significance detections () are available on the same medium-resolution grating where the [O III] line is measured, we prioritize measurements based on this grating. Otherwise, we used the grating that provides the highest S/N flux measurement. We corrected for cross-grating flux calibration offsets by using the brightest line that is covered in both the medium-resolution and prism gratings. At we avoided the [O III] and [O III] lines, since they are often blended in prism spectra. Therefore the flux calibration line is often H or H. When the medium-resolution flux measurement of a line is adopted, its flux is first normalized by the calibration line flux measured in the same grating, and then multiplied by the calibration line flux measured in the prism grating.
We also adopted the NIRSpec MSA line fluxes and EW(H) measurements for 33 galaxies from the literature with available direct-method metallicity measurements. This includes 10 galaxies from Nakajima et al. (2023), 14 galaxies from Sanders et al. (2024), and 9 galaxies from Morishita et al. (2024). We note that Laseter et al. (2024) reported 10 galaxies in the JADES DR1 data with direct-method metallicity measurements, which were independently confirmed by the pipeline detailed above.
2.2 Ground-based
Our ground-based spectra consists of 1081 galaxies selected from the archival SDSS spectra (Abazajian et al., 2009); 103 galaxies from the Nakajima et al. (2023) compilation of extremely metal-poor galaxies; 17 galaxies from the Sanders et al. (2020) compilation of direct-method metallicity measurements; and 12 galaxies from the MUSE Ultra Deep Field observations (Revalski et al., 2024). Except for the SDSS galaxies, the line fluxes and EW(H) for this sample are adopted from the corresponding papers.
We selected the SDSS galaxies from the MPA-JHU catalog (Tremonti et al., 2004; Brinchmann et al., 2004). We searched for galaxies where all of the [O II]; [O III]; H; [O III]; and H lines are detected with . We sift out the AGNs using the BPT diagram classifications of Brinchmann et al. (2004). We adopted the line fluxes and EW(H) as reported in the MPA-JHU catalog. Because the [O II] flux is not reported in any of the publicly available SDSS catalogs, we used pPXF to measure its flux for the selected galaxies. Out of the 1081 selected galaxies, 876 galaxies exhibit significant [O II] detections ().
2.3 SDSS stacks
Andrews & Martini (2013) and Curti et al. (2017) showed that the individual SDSS spectra can be stacked to enhance the [O III] signal and enable direct-method metallicity measurements for the less-explored high-metallicity () region of the parameter space. Employing a similar approach, we selected 58207 non-AGN spectra from the MPA-JHU catalog with [O II]; H; [O III]; and H detections.
We stacked these spectra on a three-dimensional grid of reddening-corrected O2, O3, and EW(H). This is in contrast with Curti et al. (2017), where the spectra are stacked on a 2-dimensional grid of reddening-corrected O2 and O3. We chose the 3-dimensional grid because our strong-method calibration relies on O2, O3, and EW(H) to estimate the gas-phase metallicity (see Section 4). We binned the O2 axis in 0.1 dex intervals, the O3 axis in 0.1 dex intervals, and the EW(H) axis in 1 dex intervals. We stacked the spectra using the stacking algorithm detailed in Langeroodi & Hjorth (2024). We used pPXF to measure the line fluxes and EW(H) for the stacks. As reported by Curti et al. (2017), the [Fe II] line is a common source of systematic offsets in [O III] flux measurement of very high metallicity galaxies. To avoid such systematics, we add the [Fe II] line to the list of emission lines fitted by pPXF. We identified 126 stacks with robust [O III] detections (), 108 of which also exhibit significant [O II] detections ().
3 Direct Measurements
We measure the ionic oxygen abundances and gas-phase metallicities (O/H) by modelling the emission lines with a 2-zone H II region (Stasińska, 1982; Garnett, 1992). This corresponds to a bithermal nebula model, where the low-ionization zone containing species such as and the high-ionization zone containing species such as are traced by different temperatures. Assuming an electron density (), temperature-sensitive line ratios can be used to calculate the electron temperature of each zone. In turn, these temperature measurements allow to derive the ionic abundances of each zone from the abundance-sensitive line fluxes. Where available, we use the [O II] and [S II] lines to estimate the electron densities (adopting the Kisielius et al. 2009 and Tayal & Zatsarinny 2010 collision strengths, respectively), while assuming otherwise. The derived temperatures and abundances are only weakly sensitive to the assumed electron density at the density regimes common for galaxies (see, e.g., Curti et al., 2017; Nakajima et al., 2023; Isobe et al., 2023). We describe the electron temperature measurements in Section 3.1 and ionic abundances and gas-phase metallicity measurements in Section 3.2. These measurements are reported in Table 1.
3.1 Electron temperatures
We measure the and electron temperatures, denoted as (O II) and (O III), respectively from the [O II]/[O II] and [O III]/[O III] flux ratios. We used the getTemDen routine of PyNeb (Luridiana et al., 2012, 2015) for this purpose, adopting the Kisielius et al. (2009) and Palay et al. (2012) collision strengths respectively for measuring the and electron temperatures444We note that the measured electron temperatures are somewhat sensitive to the adopted collision strengths table, with reported inconsistencies as high as 500K (Nicholls et al., 2013).. This results in (O III) measurements for 1510 spectra, 1004 of which also have (O II) measurements. Figure 2 shows the (O II) and (O III) temperatures for the subsample where both measurements are available.
We find that there is a clear empirical trend between (O II), (O III), O2, O3, and EW(H). For instance, the (O II)–(O III)–O3 trend is shown in Figure 2, where the data points are color-coded with their corresponding O3 measurements. Such relations are expected from the photoionization models (Izotov et al., 2006). In particular, a linear relation between (O II) and (O III) is frequently reported in the literature (Campbell et al., 1986; Garnett, 1992; Izotov et al., 2006; Pilyugin et al., 2006b, a, 2009, 2010; Curti et al., 2017), and often proposed for estimating one temperature from the other when needed (i.e., when the required lines are not covered/detected). Although Figure 2 confirms the proposed trends for average galaxies, it also shows considerable scatter around such relations.
We capture the complex relation between these parameters non-parametrically by employing a kernel density estimate (KDE; Silverman, 1986; Scott, 1992) in the 5-dimensional space of O2, O3, EW(H), (O II), and (O III). The multivariate KDE converts the multi-dimensional distribution of data into a non-parametric estimation of the probability density function (PDF). In turn, this PDF can be used to estimate the probability of specific parameter combinations. We estimate the 5-dimensional PDF using the scipy (Virtanen et al., 2020) implementation of the Scott (1992) KDE algorithm with Gaussian kernels (our multivariate probability density estimation is described in more detail in Appendix C.1; real-valued 5D images of this probability density estimate are available online in machine-readable format). We use the estimated PDF to set up an algorithm, which for each set of input O2, O3, EW(H), and (O III) estimates (O II). This particular configuration is chosen because at high redshifts it is often the case where [O II]; [O III]; H; and [O III] are detected, while [O II] is redshifted out of coverage. As such, it is often the case where measurements of O2, O3, EW(H), and (O III) are available, while (O II) cannot be directly measured.
For each set of input O2, O3, EW(H), (O III), and their uncertainties we make a 4-dimensional grid spanning the to range of each parameter in equally spaced intervals. Assuming that the and uncertainties describe half-Gaussian distributions, we assign a weight to each grid point in this 4-dimensional space. At each grid point, we calculate the probability along the (O II) axis in 10 K intervals. The resulting 1-dimensional PDFs are multiplied by the weights of the corresponding grid points and then combined to make a 1-dimensional (O II) PDF. This PDF is used to estimate the best-fit (O II) and its uncertainty as the highest-probability point and the region.
We evaluate the accuracy of our (O II) estimator through a leave-one-out cross-validation approach (additional validity tests are provided in Appendix C.2). In each iteration we take out one data point from the calibration sample, use the KDE on the remaining data points to estimate the 5-dimensional PDF, and apply the resulting (O II) estimator on the O2, O3, EW(H), and (O III) of the removed data point to estimate its (O II). Figure 3 shows the estimated (O II) vs. the directly measured values. The (O II) estimator is more accurate than 0.04 dex, defined as the absolute estimate vs. directly measured (O II) offset that contains of the estimates. As shown in Figure 3, the accuracy of our (O II) estimator declines to 0.1 dex at (O II) K, where the parameter space is sparsely sampled (see Figure 2). Figure 3 also shows that the accuracy of our (O II) estimator declines for high-redshift galaxies. This is mostly driven by the limited coverage of parameter space for these galaxies, where only 20 [O II] detections are available.
3.2 Metallicities
We calculate the ionic abundances from the [O III]/H line ratios, employing the getIonAbundance routine of PyNeb and assuming the (O III) electron temperatures calculated in Section 3.1. Similarly, the ionic abundances are calculated from the [O II]/H line ratios, assuming the (O II) electron temperatures calculated in Section 3.1. Whenever there is no (O II) measurement available, we used the (O II) estimator calibrated in Section 3.1 to estimate the (O II) based on the measured O2, O3, EW(H), and (O III). We assume that the and are the most abundant oxygen ions, and derive the oxygen abundances (gas-phase metallicity) as the sum of and ionic abundances.
| ID | redshift | RA | DEC | (O II) | (O III) | metallicity | |
| [deg] | [deg] | [mag] | [K] | [K] | |||
| 1181-00000095 | 3.91 | 189.12919839 | 62.21514056 | 0.59 | |||
| 1181-00000902 | 4.06 | 189.19327630 | 62.25372707 | 0.65 | |||
| 1181-00000910 | 4.41 | 189.11344299 | 62.25480338 | 1.23 | |||
| 1181-00000946 | 4.70 | 189.08587231 | 62.25904450 | 0.01 | |||
| 1181-00000956 | 5.42 | 189.10636574 | 62.25971421 | 0.84 | |||
| 1181-00000971 | 4.42 | 189.13093211 | 62.26199976 | 0.00 | |||
| 1181-00000988 | 6.31 | 189.16214905 | 62.26381085 | 0.72 | |||
| 1181-00001048 | 3.87 | 189.05831550 | 62.27255829 | 1.57 | |||
| 1181-00001083 | 3.80 | 189.16022669 | 62.27622688 | 0.69 | |||
| 1181-00001121 | 3.34 | 189.12982492 | 62.28116924 | 1.46 | |||
| 1181-00001129 | 7.09 | 189.17975271 | 62.28238705 | 0.07 | |||
| 1181-00001137 | 3.66 | 189.10576599 | 62.28337197 | 1.41 | |||
| 1181-00001240 | 3.33 | 189.11737086 | 62.29825830 | 0.91 | |||
| 1181-00002000 | 5.66 | 189.17594733 | 62.31153443 | 1.49 | |||
| 1181-00002864 | 3.36 | 189.14603821 | 62.25379417 | 2.02 | |||
| 1181-00002910 | 4.70 | 189.09764723 | 62.26758235 | 0.99 | |||
| 1181-00002916 | 3.66 | 189.10773892 | 62.26952483 | 2.51 | |||
| 1181-00003008 | 4.53 | 189.12051573 | 62.30316745 | 1.22 | |||
| 1181-00003982 | 7.13 | 189.10941261 | 62.23880148 | NaN | |||
| 1181-00004379 | 5.99 | 189.21938693 | 62.23824084 | 0.00 | |||
| 1181-00004550 | 3.24 | 189.19247822 | 62.23882485 | 0.92 | |||
| 1181-00006476 | 2.98 | 189.16094785 | 62.24473159 | 1.55 | |||
| 1181-00007351 | 6.05 | 189.10818294 | 62.24714628 | 0.89 | |||
| 1181-00007424 | 7.00 | 189.23290476 | 62.24738144 | 1.28 | |||
| 1181-00009104 | 6.82 | 189.24526916 | 62.25252927 | 0.57 | |||
| 1181-00010886 | 2.96 | 189.20164917 | 62.25993269 | 0.28 | |||
| 1181-00012067 | 4.06 | 189.20745025 | 62.26445323 | 1.05 | |||
| 1181-00013041 | 7.09 | 189.20377255 | 62.26842735 | 0.76 | |||
| 1181-00015529 | 3.87 | 189.21504396 | 62.27700749 | 1.83 | |||
| 1181-00016553 | 4.38 | 189.14360285 | 62.28054547 | 1.45 | |||
| 1181-00017997 | 3.32 | 189.18488493 | 62.28493708 | 0.27 | |||
| 1181-00018533 | 6.67 | 189.12121255 | 62.28640562 | 0.23 | |||
| 1181-00018536 | 6.81 | 189.15531435 | 62.28647145 | 1.01 | |||
| 1181-00019715 | 9.31 | 189.13832844 | 62.28986544 | NaN | |||
| 1181-00021747 | 3.16 | 189.16868321 | 62.23938764 | 0.78 | |||
| 1181-00022737 | 3.07 | 189.08222959 | 62.24503872 | 2.10 | |||
| 1181-00024266 | 2.96 | 189.16234694 | 62.25368410 | 0.00 | |||
| 1181-00025030 | 1.75 | 189.09503562 | 62.25682574 | 0.00 | |||
| 1181-00025351 | 3.13 | 189.17337061 | 62.23041045 | 2.34 |
4 Strong-line calibration
We use the distribution of the calibration data in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity for a non-parametric calibration of a strong-line metallicity estimator. Figure 4 shows 4 classic projections of the data, frequently used for the parametric calibration of the strong-line metallicity estimators. We adapt a method similar to that described in Section 3.1 for the non-parametric calibration. In brief, we use a kernel density estimate (KDE) in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity to estimate the probability density function (PDF) non-parametrically based on the distribution of the calibration data (our multivariate probability density estimation is described in more detail in Appendix C.1; real-valued 5D images of this probability density estimate are available online in machine-readable format). This PDF is then used to estimate the gas-phase metallicity for any combination of input O2, O3, and EW(H). The marginalization procedure is described in detail in Section 3.1. In this calibration, we only include the subsample with direct-method metallicity uncertainties lower than 0.2 dex.
We evaluate the accuracy of our strong-line metallicity estimator with a leave-one-out cross-validation approach, similar to that described in Section 3.1 (additional validity tests are provided in Appendix C.2). In each iteration, we exclude one data point from the calibration sample, calibrate the metallicity estimator on the remaining data, and use this estimator to estimate the metallicity of the excluded point based on its O2, O3, and EW(H) measurements. Figure 5 shows the strong-line gas-phase metallicity estimates vs. those measured by the direct method. Our metallicity estimations are more accurate than 0.09 dex, defined as the absolute strong-line vs. direct metallicity offset which contains of the estimates.
The accuracy of our strong-line metallicity estimator does not vary noticeably with redshift. We achieve a 0.09 dex accuracy at , 0.11 dex accuracy at , and a 0.12 dex accuracy at . We further confirm this by adding the source redshift as an extra dimension to the kernel density estimate and re-calibrating the strong-line metallicity estimator. Repeating the same leave-one-out cross-validation test as above, we achieve identical accuracies at , , and . This highlights that adding the redshift provides no additional information for estimating the gas-phase metallicities beyond what is already captured by O2, O3, and EW(H).
As shown in Figure 6, EW(H) and O32 are tightly correlated. Since O32 is widely accepted as an ionization parameter estimator (see, e.g., Kewley & Dopita, 2002; Hirschmann et al., 2023), this correlation suggests that EW(H) closely traces the ionization parameter as well. Photoionization models imply that the relationship between the strong-line ratios and gas-phase metallicity is sensitive to the ionization parameter (Nakajima et al., 2018, 2022). As such, including an ionization parameter estimator such as O32 or EW(H) in the strong-line metallicity calibration is expected to increase its accuracy. Indeed, Nakajima et al. (2022) showed that the parametric strong-line metallicity calibration can be improved, particularly at the low-metallicity end, by splitting calibration into three ionization branches as traced by the observed EW(H).
The tight correlation between the EW(H) and O32 (see Figure 6) might suggest that O2 and O3 should be sufficient to calibrate an optimal strong-line metallicity estimator; i.e., suggesting that including the EW(H) information is unnecessary. This is because EW(H) is accurately predicted from its tight correlation with O32, and the O32 information is already captured by including the O2 and O3. We test this by removing the EW(H) axis from our strong-line metallicity calibration, and using a KDE in the 3-dimensional space of O2, O3, and gas-phase metallicity. This slightly yet noticeably decreases the accuracy of our metallicity estimator to 0.17 dex. Hence, EW(H) is providing additional information beyond what is captured by O2 and O3. This seems intuitive from Figure 6, where the offset from the average EW(H)-O3 relation seems to correlate with both the gas-phase metallicity and redshift, as indicated by the color-coding in the left and right panels, respectively. Similarly, we find that including the EW(H) information slightly improves the accuracy of our (O II) estimator (Section 3.1).
Nonetheless, we make both our EW(H)-independent electron temperature and metallicity estimators publicly available. A strong-line metallicity calibration that does not rely on EW(H) measurements is often required, despite the discussed loss in accuracy when EW(H) information is not included. Most importantly, equivalent width measurements are contingent upon continuum detection, which is not always possible particularly for faint high-redshift galaxies. Moreover, equivalent width measurements based on NIRSpec MSA spectra can be subject to several systematic uncertainties as a result of either generic path-loss corrections (treating extended galaxies as point sources or uniformly extended sources) or mismatch between the emission lines and continuum profiles (e.g., Ferruit et al., 2022; D’Eugenio et al., 2024). These systematics are especially pronounced for massive intermediate-redshift galaxies, with resolved morphologies often filling the MSA shutters or even extending beyond them.
Similarly, it might be desirable to avoid explicitly using EW(H) in measuring gas-phase metallicities when investigating galaxy scaling relations such as the mass-metallicity and fundamental metallicity relations. This is because EW(H) is primarily recognized as a specific SFR (sSFR) tracer, with the EW(H) vs. ionization parameter correlation discussed above likely a secondary trend driven by the correlation between sSFR and ionization parameter (Kaasinen et al., 2018; Papovich et al., 2022; Reddy et al., 2023). As such, it might be preferred to use the EW(H)-independent estimator for measuring gas-phase metallicities to avoid introducing sSFR-driven systematic biases in the inferred scaling relations. We argue that this is not a major concern since EW(H) (and therefore sSFR by proxy) is accurately traced by O32 (as shown in Figure 6), which is included in our and most other calibrations. Therefore, we advise the user to include the EW(H) information where available. However, we note that fully exploring this and other potential systematics which affect scaling relations goes beyond the scope of this study.
5 Conclusion
We present genesis-metallicity, a non-parametric electron temperature and gas-phase metallicity estimator. This code is calibrated on a sample of 1510 [O III] detections at , compiled from the JWST/NIRSpec and ground-based observations. In particular, we report 122 new NIRSpec direct-method metallicity measurements at ; this corresponds to a fold increase in the sample size of directly-measured metallicities.
The electron temperature estimator is calibrated based on a kernel density estimate of the probability density function in the 5-dimensional space of O2, O3, EW(H), (O II), and (O III). We achieve a 0.04 dex accuracy in our (O II) estimates. The strong-line metallicity estimator is calibrated in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity. We achieve a 0.09 dex accuracy in our strong-line gas-phase metallicity estimates. Our calibration is universal, meaning that its accuracy does not depend on the target redshift.
Improved sampling of the sparsely populated regions of the emission line observables parameter space can further enhance the accuracy of our calibration. Therefore, we commit to keeping genesis-metallicity and its calibration data up-to-date in light of the upcoming data. The most recent version of genesis-metallicity and its calibration data can be found at https://github.com/langeroodi/genesis_metallicity (the calibration data will be made available upon publication). The version corresponding to this draft (1.2.0) is archived at Zenodo and can be found at Langeroodi (2025).
Acknowledgments
We greatly appreciate comments from the referee on both the manuscript and the public code, which substantially improved the presentation of the data and methodology, as well as the accessibility of the software. We are grateful to the statistics editor, whose feedback helped shape the discussion in Appendix C. We also appreciate comments from Chris Willott, which were crucial in identifying a bug in the initial version (v1.0) of the public code. This work was made possible by the public release of the reduced JWST NIRSpec MSA spectra acquired through the JADES and JOF programs. Moreover, we heavily used the latest release of the SDSS data (DR18). This work was supported by research grants (VIL16599, VIL54489) from VILLUM FONDEN.
Appendix A Emission Line Detections
Here, we present examples of our [O III] and [O II] emission line detections in JWST/NIRSpec medium-resolution spectroscopy. Figures 7, 8, 9, and 10 present [O III] detections. We note that the 3215-00265801 galaxy at is the highest redshift entry in our calibration sample, originally discovered by Laseter et al. (2024). Figures 11, 12, 13, and 14 present [O II] detections.
Appendix B Literature Parametric Calibrations
Parametric strong-line calibrations often rely on polynomial fits to projections of calibration data in line ratio vs. metallicity 2D planes. As discussed in Section 1, these calibrations risk overlooking the complexities of the higher-order parameter space. Most often, even the 2D projections cannot be fully captured by polynomials. This is evident from the large scatter of calibration data around the best-fit polynomials in 2D projection planes. This has resulted in large discrepancies between different strong-line calibrations reported in the literature. We show this in Figure 15, where some widely adopted parametric strong-line calibrations are compared with our calibration data.
Appendix C Multivariate KDE implementation
In this Section, we discuss the details of our multivariate KDE implementations presented in Sections 3.1 and 4. Our KDE bandwidth selection is described in Section C.1, while reliability tests beyond those presented in the main text are provided in Section C.2.
C.1 Bandwidths
In this work, we adopt Gaussian kernels with bandwidths selected through a cross-validation algorithm. Bandwidth selection is performed using the statsmodels library (Seabold & Perktold, 2010), adopting the maximum likelihood cross-validation approach suggested in Li & Racine (2007) which was originally proposed in Duin (1976). The selected bandwidths are then imported into our scipy implementation of KDE, described in Sections 3.1 and 4. The selected bandwidths for our (O II) electron temperature estimator (Section 3.1) in the 5-dimensional space of O2, O3, EW(H), (O II), and (O III) are reported in Table 2. Moreover, the selected bandwidths for our gas-phase metallicity estimator (Section 4) in the 4-dimensional space of O2, O3, EW(H), and gas-phase metallicity are reported in Table 3.
It is established that multivariate density estimates are far more sensitive to bandwidth selection than to the adopted kernel shape (Silverman, 1986; Scott, 1992). Here, we also explore how sensitive the accuracy of our electron temperature and metallicity estimators is to bandwidth selection. For this purpose, we adopt alternative bandwidths selected using the scipy implementation of Scott’s rule of thumb (Scott, 1992): while the selected bandwidths become generally larger by , the accuracy of both our electron temperature and gas-phase metallicity estimators remain intact.
C.2 Validation
As the dimensionality of the parameter space increases, multivariate KDEs should be used with caution because the curse of dimensionality leads to an exponential growth in the sample size required to maintain constant accuracy (Scott, 1992; Di Marzio & Lafratta, 1999; Nagler & Czado, 2015). This stems from an exponential increase in the variance term as the number of KDE covariates grows. Therefore, the addition of each new covariate must be carefully weighed against the trade-off between information gain and the variance inflation induced by the curse of dimensionality. As discussed in Sections 3.1 and 4, by removing the EW(H) covariate we construct lower-dimensional KDEs for both our electron temperature and gas-phase metallicity estimators, finding a decline in the accuracy of both. This indicates that by including the EW(H) axis we are winning the trade-off between information gain and variance inflation in both cases.
We further assess the reliability of our density estimates through bootstrap resampling. We generate 1000 bootstrap realizations of the calibration sample and recompute the KDEs for each realization. To evaluate the stability of our density estimates, we examine the variation of PDFs estimated in these realizations around their median values. The three panels in Figure 16 show the (O II) dependence of median PDF, its standard deviation, and its normalized standard deviation (defined as the standard deviation divided by the median value) for our 5-dimensional electron temperature estimator on a median O2, O3, EW(H), and (O III) slice. Similarly, the three panels in Figure 17 show the gas-phase metallicity dependence of median PDF, its standard deviation, and its normalized standard deviation for our 4-dimensional gas-phase metallicity estimator on a median O2, O3, and EW(H) slice. For both estimators the normalized scatter remains well below unity across most of the parameter range, indicating that the KDEs are stable. The only noticeable increase in normalized scatter occurs at the edges of the electron temperature and gas-phase metallicity parameter spaces, where the calibration sample provides limited coverage.
Similarly, Figure 18 shows the normalized standard deviation of the bootstrapped PDF for our 5-dimensional electron temperature estimator on a 2-dimensional plane of median O2, O3, and EW(H). Figure 19 shows the normalized standard deviation of the bootstrapped PDF for our 4-dimensional gas-phase metallicity estimator on a 2-dimensional plane of median O2 and EW(H). Similar to Figures 16 and 17, these Figures indicate that our density estimations are stable inside the region well covered by our calibration sample.
| parameter | bandwidth |
| (O2) | 0.05 |
| (O3) | 0.05 |
| (EW(H)) | 0.11 |
| (O III) [K] | 0.07 |
| (O II) [K] | 0.09 |
| parameter | bandwidth |
| (O2) | 0.05 |
| (O3) | 0.05 |
| (EW(H)) | 0.14 |
| 0.09 |